### v1

@@ Abstract
The impressive performance of ChatGPT and other foundation-model-based products in human language understanding has prompted both academia and industry to explore how these models can be tailored for specific industries and application scenarios. This process, known as the customization of domain-specific foundation models, addresses the limitations of general-purpose models, which may not fully capture the unique patterns and requirements of domain-specific data. Despite its importance, there is a notable lack of comprehensive overview papers on building domain-specific foundation models, while numerous resources exist for general-purpose models. To bridge this gap, this article provides a timely and thorough overview of the methodology for customizing domain-specific foundation models. It introduces basic concepts, outlines the general architecture, and surveys key methods for constructing domain-specific models. Furthermore, the article discusses various domains that can benefit from these specialized models and highlights the challenges ahead. Through this overview, we aim to offer valuable guidance and reference for researchers and practitioners from diverse fields to develop their own customized foundation models.
@@ 1 Introduction
ChatGPT has redefined people’s understanding of artificial intelligence with its outstanding performance. As its core technology, the Large Language Model (LLM) has become an essential tool for researchers and practitioners across various fields to improve their workflows. General-purpose foundation models are usually trained on large public datasets, enabling them to learn and address a wide range of common problems. However, these datasets cannot fully encompass all the specialized knowledge and technical details of certain specific domains. As a result, although general-purpose foundation models possess broad general knowledge, they lack the necessary depth to meet the complex needs of some specific fields [1]. Therefore, constructing domain-specific foundation models tailored to the needs of particular industries has become particularly important. Domain-specific foundation models, also known as industry-specific foundation models, are developed using data and applications specific to a particular field. Compared to general-purpose foundation models, they are trained with a large amount of domain-specific data, enabling them to more accurately understand and generate domain-specific professional content.
With the widespreading of ChatGPT-like products, the scope of the “foundation model” is gradually expanding. It is thus necessary to first make a clear definition of the foundation model discussed in this article so as to lay the foundation for the subsequent discussion on the customization of domain-specific foundation models. The foundation models mentioned in this article are neural network models that consists of at least one of the five modules in a general multi-modality foundation models (which will be detailed later). These models also exhibit the following characteristics:
  • Big data: utilizing a large volume of data covering various scenarios for model training to provide ample knowledge for the model.
  • Large parameters: the model possesses a huge number of parameters, sufficient to embed the knowledge implied by the big data into the model’s parameters.
@@
  • Generalization: the model exhibits a certain level of generalization, allowing it to perform well even in unknown data domains.
Based on the number of modalities a foundation model can process, they can be classified into single-modality foundation models and multi-modality foundation models, as shown in Table 1.
Table 1: Examples of foundation models with two different types
Types of foundation models Examples
Single-modality foundation models VGG [2], ResNet [3], GPT-1 [4], GPT-2 [5], GPT-3 [6], GPT-3.5 turbo, BERT [7], GLM [8, 9], LLaMA [10], LLaMA-2 [11], iGPT [12], LVM [13], SAM[14], BART [15], T5 [16], Time-LLM [17], UniTS [18], ST-LLM [19]
Multi-modality foundation models CoDi [20], CoDi-2 [21], Claude-3, GPT-4 [22], LLaVA [23], BriVL [24], ImageBind [25], NExT-GPT [26]
In the process of constructing domain-specific foundation models, a series of challenges will arise, especially in the stages of data acquisition and preprocessing. For example, the domain-specific data required may not be open-source or readily accessible, as it often has a high degree of confidentiality. Additionally, the modalities of domain-specific data might differ from those used for training general-purpose foundation models, making it difficult to adapt existing models to process this data. Furthermore, the environments where the domain-specific data collected may differ significantly from those for the pre-training datasets, resulting in domain-specific knowledge that pre-trained models are not familiar with. In general, constructing a domain-specific foundation model is challenging and costly, with significant implications for technical security, but it is expected to yield high economic benefits. Therefore, it is essential to thoroughly review and explore the methodologies for constructing these models, providing guideline for both researchers and practitioners.
It is noteworthy that previous review articles have predominantly focused on the development of general-purpose foundation models. Recently, while several review articles have begun to explore the domain-specific adaptation of foundation models, there is a significant gap in the literature for a comprehensive exploration of adaptation strategies that are applicable to foundation models of all modalities, extending beyond language, vision, or any single modality, to various application domains. We summarize representative surveys or review articles on foundation models in Table 2. This paper aims to provide researchers and practitioners interested in building domain-specific foundation model applications with methodological references by introducing the key methodologies therein. Also, practical examples and future directions will be discussed.
Table 2: Summary of representative surveys or review articles on foundation models
@@
Bariah et al. [44] Language Yes Telecommunications 2023
@@ 2 Preliminaries of multi-modality foundation models
In this section, we will elaborate on the preliminary technological basis for customizing domain-specific foundation models. We begin with the architecture of foundation models, detailing all functional modules. Then, from four perspectives - feature extraction, modality alignment, scaling laws, and emergent phenomena - we will explain the foundational technologies that enable each module to contribute to the high performance of foundation models.
@@ 2.1 Architecture of foundation models
According to state-of-the-art research on foundation models, it is widely considered that multi-modality foundation models can encompass all functionalities and structures of single-modality foundation models. Essentially, a single-modality foundation model implements only a subset of the functionalities of a multi-modality foundation model.
The five-module framework proposed for multi-modality foundation models in the paper [36] effectively encompasses the architecture of multi-modality foundation models with language as the central modality. However, the recent emergence of non-language foundation models, including vision foundation models [13, 14], graph foundation models [28, 29], time series foundation models and spatio-temporal (ST) foundation models [32, 33], indicates that the backbone of foundation models is expanding beyond language modalities. Therefore, we propose that the structure of multi-modality foundation models can be divided into the following five modules: Modality Encoder, Input Projector, Backbone Calculator, Output Projector, and Modality Decoder. Figure 1 illustrates the framework of a multi-modality foundation model with language as the central modality.
Figure 1: The framework of multi-modality foundation models.
For multi-modality foundation models, we define the set of all input modalities as $M$. Generally, a multi-modality foundation model has a central modality $C$. Using modality alignment techniques, the multi-modality foundation model projects all the modalities it can handle onto this central modality. Below, we define the five modules of the multi-modality foundation model and the input and output data for each module, laying down the foundation for describing the architecture of foundation models in this paper.
@@
According to the above definition, building a domain-specific foundation model entails choosing the necessary modules — some of which may be optional — and assembling a model tailored to the specific domain requirements, followed by training it with data pertinent to that domain.
2.2 Feature extraction
Feature extraction concerns extracting representative features from raw data. In the field of machine learning, particularly deep learning, feature extraction is a crucial step. As raw data often contains a significant amount of redundant and noisy information, feature extraction, which maps the data into a more information-dense feature space, enables the models to understand data structure and patterns more effectively.
In deep learning, neural networks can perform end-to-end feature extraction from raw data. However, this manner often requires large amounts of data and computational resources to ensure good performance and generalization. In each layer of neural networks, it transfers the output of previous layer to a new vector space. This structure allows for a flexible definition of the output dimensions of each layer without explicitly specifying the transformation. Leveraging these favorable characteristics, autoencoder learns an effective representation of the data by minimizing the reconstruction error between the input vector and the reconstructed vector. Autoencoder first compresses the input data into a low-dimensional feature vector and then projects them back into the original data space by a decoder, as illustrated in Figure 2. An important design principle behind the modality encoder and modality decoder metioned before, is pairing of these components as autoencoders for training. As a variant of autoencoders, variational autoencoder (VAE) aims to both minimize the reconstruction error and maximum the likelihood of the input data to learn the distribution of the compressed vectors. For example, in the image modality, VQGAN [48] is constructed in this manner. It has also been widely applied in subsequent generative models [49, 50].
@@ 2.2 Feature extraction
2.3 Modality alignment
In the workflow of a single-modality foundation model, the architecture excludes input and output projectors, reflecting the absence of a requirement for cross-modality data processing. In contrast, multi-modality foundation models must accommodate the processing of various data modalities, including both the primary and ancillary modalities. To achieve data conversion between modalities through input projectors and output projectors, the key is to apply modality alignment. The goal of modality alignment is to process the feature vectors of different modalities into a common feature space with the same dimension by using a loss function to characterize the correlation between feature vectors. Ideally, modality alignment should ensure that raw data of different modalities carrying the same semantic information are represented as the same point in the target feature space, which facilitates the realization of cross-modality information transfer.
@@ 2.3 Modality alignment
    where query $Q$, key $K$, and value $V$ are all intermediate features based on the input feature, and $d_{k}$ refers to the dimension of $Q$ and $K$. Self-attention-based mechanisms require concatenating the feature vectors of the main and auxiliary modalities as input to the transformer to generate $Q$, $K$, and $V$, allowing the model to automatically focus on the features of different modalities and achieve cross-modality information fusion. For example, VL-BERT [52] concatenates the feature vectors of text and images, and then uses the self-attention mechanism of transformer to aggregate and align language-visual features. On the other hand, methods based on cross-attention mechanism calculate $Q$, $K$, and $V$ separately for the feature vectors of the two modalities, thereby achieving cross-modality information fusion. For example, DiT [50] captures the relationship between text and images by self-attention mechanism and then realizes the image generation controlled by text. We illustrates the two fusion encoder architectures in Figure 3.
@@
    Dual encoder: As a multi-modality learning strategy, dual encoder trains a dedicated encoder for each modality independently. The core idea of this architecture is to guide the learning process of the two encoders in synchronization through semantic similarity metrics by leverage contrastive learning methods. Then the output feature vectors of different encoders can be projected into the same vector space. Specifically, this model is based on the hypothesis that if the feature vectors output by the two encoders belong to the same feature space, the feature vectors with paired labels should be closer in the vector space, and vice versa. Through this alignment method, we can expect that the output results of encoders for different modalities describing similar objects or scenes will be close enough. And even in an ideal state, they will converge to the same point in the feature space. The key to achieving this goal is to construct a reasonable model architecture and thoroughly train it on a large-scale dataset. Figure 4 shows the processing logic of the dual encoder.
@@
    However, the cost of one-to-one aligning each pair of modalities would be very high. Besides, obtaining a dataset with each pair of modalities aligned is also challenging. To this end, some researchers have proposed the bridging alignment (also known as binding alignment) strategy. This method matches all other modalities with a central modality, thereby achieving the alignment of all modalities in the semantic space. For example, ImageBind [25] takes image as the central modality, while CoDi [20] takes text as the central modality. In this way, they effectively simplify the training process of multi-modality alignment and improve the practicality and efficiency of the model. In contrastive learning, the class imbalance in long-tailed datasets can lead to reduced effectiveness, as most contrastive pairs consist of major classes instead of tail classes, making the model ineffective for the minor classes. Therefore, [53] proposed a method to solve this problem.
2.4 Scaling law
@@
Scaling law refers to the mathematical pattern about how the performance of a system changes as the scale of the system increases. In the field of AI, especially in the research and application of foundation models, scaling law describes a series of rules and phenomena about how model performance changes as the model scale expands, including the number of parameters, dataset size, computational resources. It uses quantitative analysis methods to reveal the intrinsic mechanism of the performance improvement of foundation models.
@@
[54] discussed how the inductive biases of different models affect the relationship between model scale expansion and performance. They found that model architecture is indeed one of the key factors affecting the benefits of model expansion. They also pointed out that although the standard transformer architecture may not always achieve the best performance, it does exhibit the best scalability. In the fields of computer vision [55] and natural language process [56], models based on the transformer architecture have shown an exponential relationship between model scale and model performance.
@@
Besides, [57] examined the impact of the number of downstream tasks and model scale on the performance of instruction fine-tuning. They fine-tuned the models on a wide variety of tasks by a multi-task joint training method. As a result, the language model could learn a broader language representation and knowledge, thereby enhancing its generalization ability on unseen tasks. During the joint training process, knowledge transfer between different tasks is promoted through parameter sharing, significantly improving the model’s generalization ability and performance. In addition, joint training also reduces the time and computational resources required to train each task individually, improving training efficiency. This phenomenon of model performance improving as task diversity increases is a manifestation of the scaling law. [58] verified the phenomenon of model performance increasing with the number of tasks on the large-scale benchmark OPT-IML Bench. Additionally, there are also studies that have separately provided the model performance of natural language models [56] and autoregressive generative models of various modalities [59] at different scales.
@@
Although there is no unified form for the quantitative representation of the scaling law, it can be generally represented as an exponential relationship between the model loss function and the model’s parameters, dataset size, and computational resources. We use the loss function $L(\cdot)$ to characterize the model’s performance, where a smaller loss function value represents better model performance. Eq. (7) describes the performance of a model with a given number of parameters trained to convergence on a sufficiently large dataset, where $L(N)$ is the loss function, $N$ is the number of trainable parameters of the model, $N_{c}$ is a constant, and $\alpha_{N}$ is the power law exponent. Eq. (8) provides the performance of a suitably sized model trained on a sufficiently large dataset under a given computational resource constraint, where $L(C)$ is the loss function, $C$ is the given computational resource, $C_{c}$ is a constant, and $\alpha_{C}$ is the power law exponent. Eq. (9) describes the performance of a foundation model trained with early stopping strategy on a given dataset size, where $L(D)$ is the loss function, $D$ is the dataset size (in tokens), $D_{c}$ is a constant, and $\alpha_{D}$ is the power law exponent.
@@ 2.5 Emergent phenomenon
The emergent phenomenon essentially reveals the source of the superior performance of foundation models. In the field of deep learning, especially in the domain of LLM, the emergent phenomenon has been widely observed. For example, models like LLaMA have demonstrated exceptional comprehension, generation capabilities, and even a certain degree of logical reasoning ability, which small language models cannot achieve. As the model scale increases, the model can have more parameters and a more complex structure, allowing it to capture the complex features and patterns in the data. Foundation models often exhibit strong generalization capabilities, which is mainly because their abundant parameters can store rich knowledge, enabling them to make accurate inferences and predictions on unseen data. It can provide them adaptability and universality to different tasks, and even allows the model to truly learn the underlying principles and reasoning methods in the data. [60] suggested that the emergence point of the emergent phenomenon in LLMs can be influenced by the task and the prompting method used. Specifically, the use of a chain-of-thought prompting approach can significantly enhance the LLMs’ ability to handle complex reasoning tasks [61], thereby bringing forward the emergence point of the emergent phenomenon.

### v2

@@ Abstract
The impressive performance of ChatGPT and other foundation-model-based products in human language understanding has prompted both academia and industry to explore how these models can be tailored for specific industries and application scenarios. This process, known as the customization of domain-specific foundation models (FMs), addresses the limitations of general-purpose models, which may not fully capture the unique patterns and requirements of domain-specific data. Despite its importance, there is a notable lack of comprehensive overview papers on building domain-specific FMs, while numerous resources exist for general-purpose models. To bridge this gap, this article provides a timely and thorough overview of the methodology for customizing domain-specific FMs. It introduces basic concepts, outlines the general architecture, and surveys key methods for constructing domain-specific models. Furthermore, the article discusses various domains that can benefit from these specialized models and highlights the challenges ahead. Through this overview, we aim to offer valuable guidance and reference for researchers and practitioners from diverse fields to develop their own customized FMs.
@@ 1 Introduction
Figure 1: The organization structure of this article.
ChatGPT has redefined people’s understanding of artificial intelligence with its outstanding performance. As its core technology, the Large Language Model (LLM) has become an essential tool for researchers and practitioners across various fields to improve their workflows. General-purpose foundation models (FMs) are usually trained on large public datasets, enabling them to learn and address a wide range of common problems. However, these datasets cannot fully encompass all the specialized knowledge and technical details of specific domains. As a result, although general-purpose FMs possess broad general knowledge, they lack the necessary depth to meet the complex needs of some specific fields [1]. Therefore, constructing domain-specific FMs tailored to the needs of particular industries has become particularly important. The term “domain” refers to a specific area of knowledge with logical completeness and knowledge complexity, such as healthcare, finance, or legal services. Often, only professionally trained one can master a certain domain’s terminology, concepts, and techniques. Domain-specific FMs, or industry-specific FMs, are developed using data specific to a particular field. Compared to general-purpose FMs, they are trained with a large amount of domain-specific data, enabling them to more accurately understand the unique language, terminology, concepts, and nuances inherent to the specific domain and generate domain-specific professional content.
With the widespread use of ChatGPT-like products, the scope of the “foundation model” is gradually expanding. It is necessary to clearly define the foundation model discussed to support the subsequent discussion on customizing domain-specific FMs. The FMs mentioned in this article are neural network models that consist of at least one of the five modules (which will be detailed later) in a general multi-modality FM. These models also exhibit the following characteristics:
  • Big data: The model utilizes a large volume of data covering various scenarios for training to provide ample knowledge for the model.
  • Large parameters: The model possesses a huge number of parameters sufficient to embed the knowledge implied by the big data into the model’s parameters.
@@
  • Fast adaptation capabilities: The model possesses a vast knowledge repository, enabling it to deliver robust performance even in unknown data domains with minimal fine-tuning.
In the context of artificial intelligence, we use “modality” to categorize data with a consistent data structure and semantic structure. Such a definition brings the conceptual consistency of each modality, which is the theoretical foundation of multi-modality data analysis and processing. Based on the number of modalities an FM can process, they can be classified into single-modality FMs and multi-modality FMs, as shown in Table 1. With the certain modality an FM can process, we can use different names to categorize FMs, such as large language models, large vision models, large vision-language models, etc.
Table 1: Examples of FMs with two different types
Types of FMs Examples
Single-modality FMs VGG [2], ResNet [3], GPT-1 [4], GPT-2 [5], GPT-3 [6], GPT-3.5 turbo, BERT [7], GLM [8, 9], LLaMA [10], LLaMA-2 [11], iGPT [12], LVM [13], SAM[14], BART [15], T5 [16], Time-LLM [17], UniTS [18], ST-LLM [19]
Multi-modality FMs CoDi [20], CoDi-2 [21], Claude-3, GPT-4 [22], LLaVA [23], BriVL [24], ImageBind [25], NExT-GPT [26]
In constructing domain-specific FMs, a series of challenges will arise, especially in the data acquisition and preprocessing stages. For example, the domain-specific data required may not be open-source or readily accessible, as it often has high confidentiality. Additionally, the modalities of domain-specific data might differ from those used for training general-purpose FMs, making it difficult to adapt existing models to process this data. Furthermore, the environments where the domain-specific data is collected may differ significantly from those for the pre-training datasets, resulting in domain-specific knowledge that pre-trained models are unfamiliar with. In general, constructing a domain-specific FM is challenging and costly, with significant implications for technical security, but it is expected to yield high economic benefits. Therefore, it is essential to thoroughly review and explore the methodologies for constructing these models, providing guidelines for both researchers and practitioners. The organization structure of this article is summarized in Fig. 1.
Notably, previous review articles have predominantly focused on developing general-purpose FMs. Recently, while several review articles have begun to explore the domain-specific adaptation of FMs, there is a significant gap in the literature for a comprehensive exploration of adaptation strategies that apply to FMs of all modalities, extending beyond language, vision, or any single modality, to various application domains. We summarize representative surveys or review articles on FMs in Table 2. This paper aims to provide researchers and practitioners interested in building domain-specific FM applications with methodological references by introducing the key methodologies. Also, practical examples and future directions will be discussed.
Table 2: Summary of representative surveys or review articles on FMs
@@
Bariah et al. [44] Language Yes Telecommunications 2024
@@ 2 Preliminaries of multi-modality foundation models
This section will elaborate on the preliminary technological basis for customizing domain-specific FMs. We begin with the architecture of FMs, detailing all functional modules. Then, from two perspectives - model training and scaling gain - we will explain the foundational technologies that enable each module to contribute to the high performance of FMs. Finally, we compare the performance of various general-purpose large multi-modality models for an intuitive sense of the performance that FMs can achieve and guide model employers in choosing the most suitable FM.
@@ 2.1 Architecture of foundation models
According to state-of-the-art research on FMs, it is widely considered that multi-modality FMs can encompass all functionalities and structures of single-modality FMs. Essentially, a single-modality FM implements only a subset of the functionalities of a multi-modality FM.
The five-module framework proposed for multi-modality FMs in the paper [36] effectively encompasses the architecture of multi-modality FMs with language as the central modality. However, the recent emergence of non-language FMs, including vision FMs [13, 14], graph FMs [28, 29], time series FMs and spatio-temporal (ST) FMs [32, 33], indicates that the backbone of FMs is expanding beyond language modalities. Therefore, we propose that the structure of multi-modality FMs can be divided into the following five modules: Modality Encoder, Input Projector, Backbone Calculator, Output Projector, and Modality Decoder. Figure 2 illustrates the framework of a multi-modality FM with language as the central modality.
Figure 2: The framework of multi-modality FMs with language as the central modality.
For multi-modality FMs, we define the set of all input modalities as $M$. Generally, a multi-modality FM has a central modality $C$. Using modality alignment techniques, the multi-modality FM projects all the modalities it can handle onto this central modality. Below, we define the five modules of the multi-modality FM and the input and output data for each module, laying down the foundation for describing the architecture of FMs in this paper.
@@
According to the above definition, building a domain-specific FM entails choosing the necessary modules — some of which may be optional — and assembling a model tailored to the specific domain requirements, followed by training it with data pertinent to that domain.
Here, we provide a concrete example for better illustration. CoDi-2 [21] is a language-centric multi-modality FM capable of handling various modalities such as images, videos, text, and audio. It utilizes a large language model (LLM) to generate new content. CoDi-2 employs Vision Transformers (ViTs) pre-trained by ImageBind [25] as the MEs for the image, video, and audio modalities. For the text modality, the ME is a text tokenizer. Since the output features of the ImageBind pre-trained ViT are specific to the image modality, CoDi-2 trains a multilayer perceptron (MLP) as the IP to map the features from the Image, Video, and Audio MEs to the central modality, Text. Subsequently, these Text features are fed into its BC, LLaMA2—an LLM—that generates content matching the user’s input prompt in the Text modality. Then, the generated Image and Audio content are mapped back to their respective feature spaces through an OP, another MLP. Finally, pre-trained diffusion models are used as Image, Video, and Audio MDs to generate the content.
2.2 Model training
After constructing an FM module by module, we need to train it. Model training involves setting up various training tasks to enable FM to specifically extract the features required for solving each task. Feature extraction concerns extracting representative features from raw data. As raw data often contains a significant amount of redundant and noisy information, feature extraction, which maps the data into a more information-dense feature space, enables the models to understand data structure and patterns more effectively. Feature space is the set of coordinates representing data samples across multiple feature dimensions, which mathematically refer to a high-dimensional space. Note that “raw data” and “feature” are used in this paper to differentiate whether a dataset has undergone feature extraction, which is essentially a set of mathematical calculations that increase the information density of the data. In deep learning, neural networks can perform end-to-end feature extraction from raw data. However, this often requires large amounts of data and computational resources to ensure good performance and generalization. In each layer of neural networks, it transfers the output of the previous layer to a new vector space. This structure allows for a flexible definition of the output dimensions of each layer without explicitly specifying the transformation.
@@ 2.2 Model training
Numerous training methods are available to achieve better feature extraction. They construct different loss functions to teach the model how to focus on different aspects of the data, thereby extracting various feature representations. If labeled information is available as the target for a specific task during training, supervised learning can be used to train the model’s feature extraction capabilities. For example, suppose we have labeled data for a classification task. In that case, we can construct a loss function such as Cross-Entropy Loss to measure the difference between the model’s predicted probability distribution and the true label’s probability distribution. This provides the model with a specific direction for optimization to achieve the best feature extraction for the classification task. Sometimes, we may not have labeled data for a specific task but still wish to mine features from the data to perform better on downstream tasks in the future. In such cases, unsupervised learning methods can be employed. Unsupervised training involves learning from data without the need for labeled examples. Common methods include autoencoders, which learn to compress and reconstruct input data, as shown in Fig 3, capturing important features. Another approach is using generative models like Generative Adversarial Networks (GANs) or Variational Autoencoders (VAEs) to learn the distribution of the data. Self-supervised training is also considered a subset of unsupervised training [48], such as in training LLMs where autoregressive prediction tasks are used. By generating subsequent content based on the previous content, the model inherently captures the distribution of text sequences. Combining unsupervised and self-supervised training can form semi-supervised training, which is well-suited for situations where there is a limited amount of labeled data and a large amount of unlabeled data.
Humans can also play a significant guiding role in training the model’s feature extraction capabilities. Reinforcement Learning from Human Feedback (RLHF) is a method that combines reinforcement learning with human feedback, adjusting the model based on human evaluations of its behavior to learn feature representations that align more closely with human expectations. This technique is often mentioned in the context of training LLMs. First, the model generates a series of outputs and requests a human evaluation of these outputs, such as scoring, ranking, or providing other forms of feedback. These feedback data are then used to train a reward model that assesses the quality of different outputs based on human preferences. Then, using the reward model, the initial model is fine-tuned through reinforcement learning algorithms. During this fine-tuning process, the model continuously adjusts its behavior strategy based on the feedback from the reward model to maximize cumulative rewards. Through this process, the model can better understand user preferences and needs, generating outputs more aligned with human expectations.
For a single-modality FM, cross-modality data is unavailable, so the architecture excludes IPs and OPs. In contrast, multi-modality FMs must accommodate the processing of various data modalities, including the primary and ancillary modalities. To achieve data conversion between modalities through IPs and OPs, the key is to apply modality alignment. The goal of modality alignment is to process the feature vectors of different modalities into a common feature space with the same dimension by using a loss function to characterize the correlation between feature vectors. Ideally, modality alignment should ensure that raw data of different modalities carrying the same semantic information are represented as the same point in the target feature space, facilitating cross-modality information transfer.
@@
    where query $Q$, key $K$, and value $V$ are all intermediate features based on the input feature, and $d_{k}$ refers to the dimension of $Q$ and $K$.
@@
    Self-attention-based mechanisms require concatenating the feature vectors of the main and auxiliary modalities as input to the transformer to generate $Q$, $K$, and $V$, allowing the model to automatically focus on the features of different modalities and achieve cross-modality information fusion. For example, VL-BERT [50] concatenates the feature vectors of text and images, and then uses the self-attention mechanism of transformer to aggregate and align language-visual features. On the other hand, methods based on the cross-attention mechanism calculate $Q$, $K$, and $V$ separately for the feature vectors of the two modalities, thereby achieving cross-modality information fusion. For example, DiT [51] captures the relationship between text and images by self-attention mechanism and then realizes the image generation controlled by text. We illustrate the two fusion encoder architectures in Figure 4.
@@
    Dual encoder: As a multi-modality learning strategy, dual encoder trains a dedicated encoder for each modality independently. The core idea of this architecture is to guide the learning process of the two encoders in synchronization through semantic similarity metrics by leveraging contrastive learning methods. Then, the output feature vectors of different encoders can be projected into the same vector space. Specifically, this model is based on the hypothesis that if the feature vectors output by the two encoders belong to the same feature space, the feature vectors with paired labels should be closer in the vector space and vice versa. Through this alignment method, we can expect that the output results of encoders for different modalities describing similar objects or scenes will be close enough. And even in an ideal state, they will converge to the same point in the feature space. The key to achieving this goal is to construct a reasonable model architecture and thoroughly train it on a large-scale dataset. Figure 5 shows the processing logic of the dual encoder.
@@
    However, the cost of one-to-one aligning each pair of modalities would be very high. Besides, obtaining a dataset with each pair of modalities aligned is also challenging. To this end, some researchers have proposed the bridging alignment (also known as binding alignment) strategy. This method matches all other modalities with a central modality, thereby achieving the alignment of all modalities in the semantic space. For example, ImageBind [25] takes the image as the central modality, while CoDi [20] takes text as the central modality. In this way, they effectively simplify the training process of multi-modality alignment and improve the practicality and efficiency of the model. In contrastive learning, the class imbalance in long-tailed datasets can lead to reduced effectiveness, as most contrastive pairs consist of major classes instead of tail classes, making the model ineffective for the minor classes. Therefore, [52] proposed a method to solve this problem.
2.3 Gain of scaling up
As the training scale of artificial intelligence (AI) models continues to grow, their capabilities in various tasks demonstrated significant leaps in performance. When the model size reaches the FM level, researchers have introduced the concepts of scaling laws and emergent phenomena to illustrate the quantitative and qualitative relationships between training scale and model performance.
@@ 2.3 Gain of scaling up
Scaling law refers to the mathematical pattern of how the performance of a system changes as the scale of the system increases. In AI, especially in the research and application of FMs, scaling law describes rules and phenomena about how model performance changes as the model scale expands, including the number of parameters, dataset size, and computational resources. It uses quantitative analysis methods to reveal the intrinsic mechanism of the performance improvement of FMs.
@@
[53] discussed how the inductive biases of different models affect the relationship between model scale expansion and performance. They found that model architecture is indeed one of the key factors affecting the benefits of model expansion. They also pointed out that although the standard transformer architecture may not always achieve the best performance, it does exhibit the best scalability. In computer vision [54] and natural language process [55], models based on the transformer architecture have shown an exponential relationship between model scale and model performance.
@@
Besides, [56] examined the impact of the number of downstream tasks and model scale on the performance of instruction fine-tuning. They fine-tuned the models on various tasks by a multi-task joint training method. As a result, the language model could learn a broader language representation and knowledge, enhancing its generalization ability on unseen tasks. During the joint training process, knowledge transfer between different tasks is promoted through parameter sharing, significantly improving the model’s generalization ability and performance. In addition, joint training also reduces the time and computational resources required to train each task individually, improving training efficiency. This phenomenon of model performance improving as task diversity increases is a manifestation of the scaling law. [57] verified the phenomenon of model performance increasing with the number of tasks on the large-scale benchmark OPT-IML Bench. Additionally, some studies have separately provided the model performance of natural language models [55] and autoregressive generative models of various modalities [58] at different scales.
@@
Although there is no unified form for the quantitative representation of the scaling law, it can be generally represented as an exponential relationship between the model loss function and the model’s parameters, dataset size, and computational resources. We use the loss function $L(\cdot)$ to characterize the model’s performance, where a smaller loss function value represents better model performance. Eq. (7) describes the performance of a model with a given number of parameters trained to convergence on a sufficiently large dataset, where $L(N)$ is the loss function, $N$ is the number of trainable parameters of the model, $N_{c}$ is a constant, and $\alpha_{N}$ is the power law exponent. Eq. (8) provides the performance of a suitably sized model trained on a sufficiently large dataset under a given computational resource constraint, where $L(C)$ is the loss function, $C$ is the given computational resource, $C_{c}$ is a constant, and $\alpha_{C}$ is the power law exponent. Eq. (9) describes the performance of an FM trained with an early stopping strategy on a given dataset size, where $L(D)$ is the loss function, $D$ is the dataset size (in tokens), $D_{c}$ is a constant, and $\alpha_{D}$ is the power law exponent.
@@
The emergent phenomenon essentially reveals the source of the superior performance of FMs. In deep learning, especially in the domain of LLM, the emergent phenomenon has been widely observed. For example, models like LLaMA have demonstrated exceptional comprehension, generation capabilities, and even a certain degree of logical reasoning ability, which small language models cannot achieve. As the model scale increases, it can have more parameters and a more complex structure, allowing it to capture the complex features and patterns in the data. FMs often exhibit strong generalization capabilities because their abundant parameters can store rich knowledge, enabling them to make accurate inferences and predictions on unseen data. It can provide them adaptability and universality to different tasks and even allows the model to truly learn the underlying principles and reasoning methods in the data. [59] suggested that the task and the prompting method used can influence the emergence point of the emergent phenomenon in LLMs. Specifically, using a chain-of-thought prompting approach can significantly enhance the LLMs’ ability to handle complex reasoning tasks [60], thereby bringing forward the emergence point of the emergent phenomenon.